# Real-World Adversarial Attacks on RF-Based Drone Detectors

### v1

@@ Abstract
Index Terms—  Drone detection, Adversarial ML.
@@ 1 Introduction
With unmanned aerial vehicles (UAVs) increasingly used in civilian and defense applications [1], reliable drone detection near critical infrastructure (e.g., airports, power plants) is essential for the mitigation of risks from unauthorized drone activity [1, 2, 3]. Radio frequency (RF) based detection, which can leverage drones’ spectral–temporal signatures, has shown robustness under non-line-of-sight conditions [4, 2]. Converting complex baseband (I/Q) signals to time frequency images (i.e., spectrograms) via Short-Time Fourier Transform (STFT) [1] enables the use of computer vision object detection (OD) frameworks like YOLO [5] and Faster R-CNN [6] for drone identification and localization in multi-emitter environments [1, 7, 8]. However, OD frameworks are vulnerable to adversarial attacks where small perturbations can impede detection and induce false negatives [9].
@@
We introduce the first physical adversarial attack targeting drone transmission specifically designed for RF based OD models. Rather than optimizing perturbations in the digital feature space, our approach generates adversarial waveforms directly in the I/Q domain and emits them over-the-air (OTA) from an independent transmitter. This creates interference that is combined with legitimate drone signals but eliminates the need for the complex digital-to-RF conversion and ensures compatibility with standard RF hardware. We perform both digital and physical evaluation of our attack’s robustness across various OD architectures, demonstrating the generalizability of I/Q domain adversarial perturbation in RF detection systems.
We summarize the contributions of our work as follows: (1) To the best of our knowledge, this is the first work to demonstrate a physical adversarial attack specifically targeting image based OD models, addressing the unique challenges of multi-emitter detection scenarios. (2) We introduce a novel method that optimizes perturbations directly in the I/Q domain, ensuring physical realizability and compatibility with standard RF transmission chains. (3) We conduct comprehensive OTA experiments using RF transmissions from four drone platforms, validating that modest, structured I/Q perturbations reliably degrade detection performance while overcoming real-world wireless challenges such as timing misalignment and channel effects.
@@ 2 Related Work
Physical layer validation and practical deployment. Only two studies conducted physical experiments, performing a per-sample attack [10] and UAP on an SM classifier [18, 20], leaving OTA validation for RF FP-ID tasks, especially on OD models, unexplored. Additionally, most existing attacks assume perfect knowledge of target architectures, with limited exploration of transferability [17, 15]. Critically, CUAPs have not been studied in realistic, asynchronous multi-emitter conditions where selective targeting of specific drone classes must preserve detection of legitimate traffic.
@@ 3.2 Optimization Objective
Perturbation $\delta$ is optimized through an iterative procedure using standard gradient descent, aiming to achieve our main objective of reducing target class confidence while preserving non-target classes as follows:
@@ 4.1 Dataset
[…]
@@
Attack set generation. We assume an adversary that aims to prevent their drone’s detection by OD systems. They can control their drone’s transmit channel (i.e., center frequency) but have no control of other drones’ transmission channels. Accordingly, we recorded the target drone OTA in a realistic, noisy environment; other drones’ signals were digitally mixed into the I/Q stream at random channels across the 2.4 GHz ISM band. This setup reflects a realistic threat model in which the adversary can manipulate the target drone’s emissions but cannot influence coexisting signals. The attack set used to train the adversarial waveform was split into 70%, 15%, and 15% for the training, validation, and test subsets, respectively.
@@ 4.3 Evaluation Metrics
For digital evaluations, we report average precision (AP) for the target class and mean AP (mAP) across all non-target classes, where AP represents the area under the precision–recall curve at an IoU threshold of 0.5. For physical experiments, we quantify performance using the missed detection rate (MDR) for the class, where non-target classes are aggregated in a single class:
@@ 4.4 Experimental Results
Table 1: The AP (for SPR = 10 dB) in the following settings: clean, AWGN, WB, GB, CS, and $B$.
@@
Physical attacks. Due to space limitations, we only present the closed-set CUAP (shown to degrade all tested architectures in digital evaluations) performance in a realistic multi-emitter environment to demonstrate OTA feasibility. Following the described threat model, the target drone operated in the same predetermined channel, while the other three drones independently selected transmission channels. This setup reduces temporal variability (no spectral hopping) and isolates OTA effects. The optimized $\delta$ was transmitted along with the legitimate communications, with transmit power adjusted to match the received perturbation levels from the digital experiments. All drones were positioned 3 m from the receiving antenna, consistent with the original recordings. Since OTA transmissions timing and phase are non-repeatable, Clean and Attacked settings use separate recordings acquired consecutively under identical setup.
The results presented in Table 2 highlight the transferability of our optimization approach to physical attacks. The target class MDR increased substantially, with four models achieving near-complete suppression ($\geq 90.3\%$) and YOLOv9 reaching 71.5%. Detection of the non-target classes remained largely preserved (MDR degradation $\leq 3.2\%$), confirming a class-specific targeting effect. The OTA results demonstrate that I/Q domain perturbations retain adversarial effects after propagation through physical RF channels.
@@
Setting YOLOv5 YOLOv8 YOLOv9 YOLOv11 FRCNN
Clean 1.7 / 3.6 1.0 / 2.8 0.7 / 2.3 0.4 / 5.2 3.9 / 14.7
Attacked 99.9 / 6.8 99.9 / 2.9 71.5 / 3.4 99.9 / 5.1 90.3 / 6.9
Table 2: CUAP’s performance in OTA settings. High MDR for the target class indicates a successful attack.
5 Conclusion
This work presents the first physical adversarial attack against image based RF drone detectors by learning CUAPs in the I/Q domain, enabling OTA implementation. Our evaluation highlights the attack’s selective suppression of target class detection with minimal impact on the detection of non-target classes and strong cross-architecture transfer. Our findings reveal fundamental vulnerabilities in existing detection systems and suggest the need for robust defenses for deployed RF based OD models.

### v2

@@ Abstract
Index Terms—  Drone detection, Adversarial ML., Physical Attack, RF Sensing, Over-the-Air Attack, I/Q domain
@@ 1 Introduction
With unmanned aerial vehicles (UAVs) increasingly used in civilian and defense applications [1], reliable drone detection near critical infrastructure (e.g., airports, power plants) is essential for the mitigation of risks from unauthorized drone activity [1, 2]Radio frequency (RF) based detection, which can leverage drones’ spectral–temporal signatures, has shown robustness under non-line-of-sight conditions [3, 2]. Converting complex baseband (I/Q) signals to time frequency images (i.e., spectrograms) via Short-Time Fourier Transform (STFT) [1] enables the use of computer vision object detection (OD) frameworks like YOLO [4] and Faster R-CNN [5] for drone identification and localization in multi-emitter environments [1, 6, 7]. However, OD frameworks are vulnerable to adversarial attacks where small perturbations can impede detection and induce false negatives [8].
@@
We introduce the first physical adversarial attack targeting drone transmission specifically designed for RF based OD models. Rather than optimizing perturbations in the digital feature space, our approach generates adversarial waveforms directly in the I/Q domain and emits them over-the-air (OTA) from an independent transmitter. This creates interference that is combined with legitimate drone signals but eliminates the need for the complex digital-to-RF conversion and ensures compatibility with standard RF hardware. We perform both digital and physical evaluation using the same optimized I/Q-domain CUAP across various OD architectures across various OD architectures, demonstrating the generalizability of I/Q domain adversarial perturbation in RF detection systems.
We summarize the contributions of our work as follows: (1) To the best of our knowledge, this is the first work to demonstrate a physical adversarial attack specifically targeting image based OD models, addressing the unique challenges of multi-emitter detection scenarios. (2) We introduce a novel method that optimizes perturbations directly in the I/Q domain, ensuring physical realizability and compatibility with standard RF transmission chains. (3) We conduct comprehensive OTA experiments using RF transmissions from four drone platforms, validating that modest, structured I/Q perturbations learned digitally reliably degrade detection performance while overcoming real-world wireless challenges such as timing misalignment and channel effects.
@@ 2 Related Work
Physical layer validation and practical deployment. Only a few studies conducted physical experiments, performing a per-sample attack [9] and UAP on an SM classifier [17, 19], leaving OTA validation for RF FP-ID tasks, especially on OD models, unexplored. Additionally, most existing attacks assume perfect knowledge of target architectures, with limited exploration of transferability [16, 14]. Critically, CUAPs have not been studied in realistic, asynchronous multi-emitter conditions where selective targeting of specific drone classes must preserve detection of legitimate traffic.
@@ 3.2 Optimization Objective
Perturbation $\delta$ is optimized digitally through an iterative procedure using standard gradient descent, and then reused unchanged for both digital and OTA evaluations, aiming to achieve our main objective of reducing target class confidence while preserving non-target classes as follows:
@@ 4.1 Dataset
Training data collection. We collected RF baseband I/Q data using an Analog Devices ADRV9009 transceiver interfaced to a Xilinx ZCU102 evaluation board. All training recordings were obtained in a shielded room (Faraday cage) to suppress external interference and ensure signal integrity. The receiver was configured with a local oscillator at 2.45 GHz, a complex sampling rate of 122.88 $\mathrm{MS/s}$, and an effective analog bandwidth of 100 MHz, fully covering the 2.4 GHz ISM band used by our drones. Four types of drones (Mavic 2 Zoom, Mavic Pro, Mavic Air, and Matrice 600) were operated in idle-on mode separately on multiple RF channels; each drone’s transmissions were recorded, converted to spectrograms using the pipeline described in Sec. 3.1, and segmented into detector input frames of size $F\times T=1,024\times 1,024$ (frequency $\times$ time).
@@
Attack set generation. We assume an adversary that aims to prevent their drone’s detection by OD systems. They can control their drone’s transmit channel (i.e., center frequency) but have no control of other drones’ transmission channels. Accordingly, we recorded the target drone OTA in a realistic, noisy environment; other drones’ signals were digitally mixed into the I/Q stream at random channels across the 2.4 GHz ISM band. This setup reflects a realistic threat model in which the adversary can manipulate the target drone’s emissions via a nearby adversarial transmitter but cannot influence coexisting signals. The attack set used to train the adversarial waveform was split into 70%, 15%, and 15% for the training, validation, and test subsets, respectively.
@@ 4.3 Evaluation Metrics
For digital evaluations, we report average precision (AP) for the target class and mean AP (mAP) across all non-target classes, where AP represents the area under the precision–recall curve at an IoU threshold of 0.5. For physical experiments, where precise frame-level annotations are unavailable, we quantify performance using the missed detection rate (MDR) for the class, reflecting the same attack objective as AP, where non-target classes are aggregated in a single class:
@@ 4.4 Experimental Results
Table 1: The AP (for SPR = 10 dB) in the following settings: clean, AWGN, WB, GB, CS, and $B$. $\downarrow$ indicates that lower values are better and $\uparrow$ indicates that higher values are better.
@@
Physical attacks. Due to space limitations, we present only the closed-set CUAP, which exhibited the strongest cross-architecture transfer in digital evaluations and represents a practical scenario when the target detector architecture is unknown, demonstrating OTA feasibility in a realistic multi-emitter environment. Drones were positioned at fixed distances (3m, 5m, and 7m) from the receiving antenna to assess range effects. Since OTA transmissions timing and phase are non-repeatable, Clean and Attacked settings use separate recordings acquired consecutively under identical setup.
The results presented in Table 2 highlight the transferability of our digital optimization approach to physical attacks. Under attack, the target class MDR reached near-complete suppression ($\geq 99\%$) on most models across distances, with YOLOv9 at 3m and FRCNN as exceptions. Detection of the non-target classes remained largely preserved (MDR degradation $\leq 18.3\%$), confirming a class-specific targeting effect. These OTA results demonstrate that I/Q domain perturbations retain adversarial effects after propagation through physical RF channels across varying ranges.
@@
Distance Setting YOLOv5 YOLOv8 YOLOv9 YOLOv11 FRCNN
3m Clean 1.7 / 3.6 1.0 / 2.8 0.7 / 2.3 0.4 / 5.2 3.9 / 14.7
Attacked 100 / 6.8 100 / 2.9 71.5 / 3.4 100 / 5.1 90.3 / 6.9
5m Clean 0.9 / 3.3 0.2 / 3.6 5.2 / 1.7 1.7 / 3.9 2.9 / 9.4
Attacked 100 / 6.5 100 / 5.8 100 / 12.3 100 / 16.0 90.7 / 3.7
7m Clean 3.0 / 2.1 1.3 / 11.0 9.6 / 1.1 3.0 / 12.3 7.6 / 11.0
Attacked 100 / 6.7 99.5 / 13.6 100 / 5.7 100 / 18.3 87.5 / 3.3
Table 2: CUAP (CS) OTA performance at different distances. High MDR for the target class indicates a successful attack. $\downarrow$ indicates that lower values are better and $\uparrow$ indicates that higher values are better.
5 Discussion and Conclusion
This work presents the first physical adversarial attack against image-based RF drone detectors by learning CUAPs in the I/Q domain, enabling OTA implementation. Our evaluation demonstrates selective suppression of the target class with minimal impact on non-target classes and strong cross-architecture transfer. These proof-of-concept results, limited to an office environment, idle-on drones, and fixed distances, reveal fundamental vulnerabilities in RF-based detection systems and motivate robust defenses such as adversarial training and RF anomaly detection.